A card is drawn and a die is thrown. Find the following probabilities:
a) Drawing a spade and rolling a four
b) Rolling an odd number and drawing an odd numbered card
c) Rolling a factor of 12 and drawing a card that is a factor of twelve (exclude 1 in both cases)

Answer by jim_thompson5910(35256)

a)

P(Spade and rolling a 4) = P(Spade)*P(rolling a 4) = (13/52)*(1/6) = 13/312

Note: There are 13 spade cards out of a total of 52. So P(Spade) = 13/52. Also, I'm assuming that we're rolling a 6 sided die, which means that P(Rolling a 4) = 1/6



b)

P(Rolling an odd number and drawing an odd numbered card) = P(Rolling an odd number) * P(Drawing an odd numbered card) = (3/6)*(16/52) = 48/312 = 2/13


Note: There 3 odd numbers on a die (1, 3, 5) out of 6. So P(Rolling an odd number) = 3/6

Also, the odd numbered cards are: 3, 5, 7, 9 and there are 4 suits. So there are 4*4 = 16 odd cards out a total of 52. So P(Drawing an odd numbered card) = 16/52


c)

I'll let you tackle this one. Let me know if you need help on it.